# Hybrid d/p-wave altermagnetism in Ca3Ru2O7 and strain-controlled spin splitting

### v2

@@ Results
Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in supplementary information file (SI) table S1). Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively. Next, we will focus on the AM phases of CRO, primarily in Configuration B.
In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table S1).
@@
We now examine the electronic structure of both configurations. Figure 3 shows the total and orbital-resolved density of states. While configuration A is metallic, configuration B exhibits a narrow gap at the Fermi level. In an ideal octahedral environment, the Ru , , and orbitals are degenerate. Orthorhombic distortions lift this degeneracy slightly, but it is the Néel-type magnetic order in the -plane (conf. B) that drives orbital differentiation by enhancing -type hybridization between Ru- and O- orbitals. This leads to a strong bonding–antibonding splitting, pushing the bonding states to lower energy and making them nearly filled. The resulting spectral redistribution lowers the chemical potential relative to the less-affected, half-filled and orbitals. The gap thus reflects an orbital-selective mechanism driven by the interplay between magnetism and crystal-field effects, as proposed for Ca2RuO4 [40]
@@
The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals. In contrast, the bands between eV, dominated by the orbital character, show negligible spin splitting (see, Fig. S3). Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [40]. When the or coordinate is inverted in reciprocal space, the non-relativistic spin splitting changes sign, while inversion of leaves it unchanged. Notably, this altermagnetic order matches that found in Ca2RuO4, despite its crystallization in a different space group [40]. The orbital-selective magnetism depends on the magnetic space group and orbital character. In conf.B, the nearly filled dxy orbitals do not contribute to magnetism, leading to orbital-selective AM. Consequently, conf. C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig. S3.
@@ Discussion
From the energy-strain relationship shown in Fig. 8(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain (). Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis. This expansion modifies the balance of structural distortions and stabilizes configuration B over configuration A. At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting. Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (Fig. S5). This effect is reflected in the rising AMQ values in Fig. 8(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.
@@ Methods
Data availability
All data supporting the findings of this study are included in the article and its Supplementary Information.
Code Availability
Density functional theory calculations were performed using the Vienna ab initio Simulation Package (VASP), accessed under an institutional license through the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2. Custom Python scripts developed for post-processing and analysis are available from the corresponding author upon reasonable request.
@@ Acknowledgments
A.L. thanks the Technische Universität Dresden Professor Fellowship Program. C. A. acknowledges the useful discussion with G. Cuono and A. Fakhredine about altermagnetism in Ca2RuO4. C. A. is supported by the Foundation for Polish Science project “MagTop” no. FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021-2027 (FENG). J.W.G. acknowledges financial support from Chilean ANID-FONDECY grants N. 1220700 and 1221301. T.B. would like to thank the German Science Foundation (DFG) for supporting this work via the German-Israeli Project Cooperation with the Grant No. HE 3543/42-1. The authors gratefully acknowledge the computing time made available to them on the high-performance computer at the NHR Center of TU Dresden and on the high-performance computers Noctua 2 at the NHR Center PC2.

### v3

@@ Results
Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in SM table S1). Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively. Next, we will focus on the AM phases of CRO, primarily in Configuration B.
| E | E-State | M-Phase | () | |
| A | 0 | Metallic | AFM | 1.44 |
| B | 34.90 | N-Insulator | AM | 1.35 |
| C | 36.46 | Metallic | AM | 1.38/1.45 |
| D | 39.22 | N-Insulator | AFM | 1.35 |
In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table 1).
@@
We now examine the electronic structure of both configurations. Figure 3 shows the total and orbital-resolved density of states. While configuration A is metallic, configuration B exhibits a narrow gap at the Fermi level. In an ideal octahedral environment, the Ru , , and orbitals are degenerate. Orthorhombic distortions lift this degeneracy slightly, but it is the Néel-type magnetic order in the -plane (conf. B) that drives orbital differentiation by enhancing -type hybridization between Ru- and O- orbitals. This leads to a strong bonding–antibonding splitting, pushing the bonding states to lower energy and making them nearly filled. The resulting spectral redistribution lowers the chemical potential relative to the less-affected, half-filled and orbitals. The gap thus reflects an orbital-selective mechanism driven by the interplay between magnetism and crystal-field effects, as proposed for Ca2RuO4 [39].
@@
The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals. In contrast, the bands between eV, dominated by the orbital character, show negligible spin splitting (see, Fig. LABEL:fig:S2). Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [39]. When the or coordinate is inverted in reciprocal space, the non-relativistic spin splitting changes sign, while inversion of leaves it unchanged. Notably, this altermagnetic order matches that found in Ca2RuO4, despite its crystallization in a different space group [39]. The orbital-selective magnetism depends on the magnetic space group and orbital character. In conf.B, the nearly filled dxy orbitals do not contribute to magnetism, leading to orbital-selective AM. Consequently, conf. C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig. LABEL:fig:S2.
@@
To further dissect the symmetry of the spin splitting in three dimensions (details provided in the SM, section SVI), we perform a multipolar analysis of the energy texture . In the non-relativistic limit, the response is purely even in , confirming the expected pure -wave altermagnetic state dictated by symmetry. Including SOC activates odd-parity components and transforms the system into a hybrid state with concurrent -, -, and -wave character. Using spinor band pairs near the Fermi level, we find overall -wave dominance, with fractional -norm weights , , and . The hybridization is strongly anisotropic across spin channels: the and components are -dominated (, ), whereas the component is -dominated ().
@@ Discussion
From the energy-strain relationship shown in Fig. 8(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain (). Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis. This expansion modifies the balance of structural distortions and stabilizes configuration B over configuration A. At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting. Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (Fig. S8 SM). This effect is reflected in the rising AMQ values in Fig. 8(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.
@@ Acknowledgments
A.L. thanks the Technische Universität Dresden Professor Fellowship Program. C. A. acknowledges the useful discussion with G. Cuono and A. Fakhredine about altermagnetism in Ca2RuO4. C. A. is supported by the Foundation for Polish Science project “MagTop” no. FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021-2027 (FENG). J.W.G. acknowledges financial support from Chilean ANID-FONDECY grants N. 1220700 and 1221301. T.B. would like to thank the German Science Foundation (DFG) for supporting this work via the German-Israeli Project Cooperation with the Grant No. HE 3543/42-1.
@@ 1 Competing interests
Supplementary Information
SI Ca3Ru2O7 structural parameters
Table S1 shows the energy difference regarding the ground state system (Configuration A) considering spin orbit coupling (SOC) and without SOC (NSOC), respectively (considering the spins along direction). , and are the lattice parameter (obtained from NSOC calculations).
| Configuration | ENSOC (meV/Ru) | ESOC (meV/Ru) | a/b (Å) | c (Å) |
| A | 0 | 0 | 5.415/5.601 | 19.54 |
| B | 34.90 | 36.86 | 5.415/5.726 | 19.23 |
| C | 36.46 | 60.55 | 5.398/5.652 | 19.449 |
| D | 39.22 | 37.32 | 5.419/5.722 | 19.261 |
| Exp [1] | – | – | 5.367/5.535 | 19.521 |
Note: The calculations including spin-orbit coupling were performed using the relaxed structures obtained from collinear calculations and do not include further structural relaxation.
Effect Hubbard-U energy differences: Table S2 shows the energy difference relative to the ground state (Configuration A) for three different values of the Hubbard parameter. We observe that the choice of does not significantly affect the relative stability of the magnetic phases with respect to configuration A. For eV, configuration A begins to exhibit structural changes, with distortions in the octahedral environment indicating the onset of structural instability. This behavior has been previously investigated [2], and it has been established that at eV the system stabilizes in a different structural phase [3, 2].
| Configuration | E (U = 0.5) | E (U = 1.0) | E (U = 1.1) |
| A | 0 | 0 | 0 |
| B | 48.41 | 34.90 | 30.76 |
| C | 52.35 | 36.46 | 32.79 |
| D | 74.0 | 39.23 | 37.29 |
SII Density of states with and without Hubbard-U interaction.
Figure S1 shows the total density of states (DOS) for configurations A and B, with and without the inclusion of the Hubbard-U interaction. At U=0 eV, both systems exhibit a metallic state. When the Hubbard-
U interaction is introduced, configuration B favors a narrow-gap insulating state, whereas configuration A remains metallic.
This different response to the Hubbard-U
interaction arises from the localization versus delocalization effects associated with the symmetry of the wave functions, which in turn are determined by the magnetic configurations. Specifically, configuration B corresponds to a singlet state, where the spin part of the wave function is antisymmetric under exchange, leading to a symmetric (bonding-like) spatial part that favors electron delocalization. In contrast, configuration A resembles a triplet state, with a symmetric spin part and thus an antisymmetric (antibonding-like) spatial part, resulting in more localized electron clouds.
In essence, the Hubbard- term promotes electron localization. Since configuration A already exhibits a relatively localized electronic character, the effect of a small value is minimal. However, by increasing , a gap opens, leading to an insulating state with a narrow gap, similar to that observed in configuration B. More details about the interplay among the different degrees of freedom in configuration A can be found in Ref. [3]. In contrast, configuration B initially favors a more delocalized state; therefore, the introduction of induces localization, which perturbs the spatial part of the wave function and, consequently, alters the magnetic order.
SIII Ca3Ru2O7 altermagnetic configurations
Fig. S2 (upper panel) shows the band structure for the AM configurations of Ca3Ru2O7: configuration B (a) and configuration C (b).
Fig. LABEL:fig:S2 shows the band structure projected onto the -orbitals by occupation state for the B and C configurations. While the configuration B hosts a P-2 dyz-wave altermagnetism, the configuration C hosts a dxz-wave altermagnetism. Configuration C exhibits altermagnetism in both dxy and
dxz-dyz bands, therefore, the altermagnetism of the configuration C is not orbital selective. In the case of the B configuration, the main contributions to the AM bands are between -1.5 and -1.0 eV.
SIV Relativistic bands
Figure LABEL:fig:S4 shows the band structure for configuration B.
Upper panel: Band structure along the path, where altermagnetic splitting is most pronounced (see Fig. 4 in the main text). The projection retains strong AM features at lower energies, consistent with the collinear case. In contrast, the projection reveals AM features between eV and eV.
Lower panel: Band structure along the path. The and components retain the antiferromagnetic character, while the projection shows clear AM features near the Fermi level.
SV Band structure under strain
Figure LABEL:fig:S4 shows the variation of the DOS at the Fermi level as a function of strain. It can be observed that under compressive strain, the DOS becomes broader, while under tensile strain, it becomes more localized.
Figure S8 presents the band structure under different strain values (-plane) for two cases. The upper panels correspond to fully relaxed structures, while the lower panels depict the same strain conditions (same lattice parameters) but with atomic positions fixed at their zero-strain values (non-relaxed calculations). In this scenario, the systems retain the Ru-O rotation from the zero-strain configuration.
Notably, removing the Ru-O distortions primarily affects the band gap.
The main text focused on the absolute energy difference between majority and minority bands. However, in Fig. S10, we instead plot the signed difference (spin-up minus spin-down) to highlight how the splitting changes sign across the Brillouin zone. For instance, the splitting inverts going from to and from to . Such sign reversals showcase the complementary nature of the -dependent spin splitting that emerges in the altermagnetic state.
While taking the absolute value is helpful to quantify the splitting strength, plotting the raw sign clarifies how the splitting evolves—and sometimes flips—across distinct points in -space.
Figure S10 shows the resulting along the entire -path, defined by
where is the number of bands at each -point within the specified energy range, and is the number of occupied bands. These plots show that is sensitive to the applied strain. Under compressive strain, the overall magnitude of decreases compared to the unstrained system. In contrast, under tensile strain, it increases, underscoring how the electronic structure and spin splitting respond to lattice deformations.
SVI 3D Multipolar Analysis of the Spin-Resolved Splitting
Conventions and Brillouin-zone averaging
Let be the matrix of direct-lattice vectors (in Å) and the reciprocal-lattice matrix (in Å-1). Fractional momenta are mapped to Cartesian reciprocal coordinates by . Brillouin-zone (BZ) averages are computed discretely using the -point integration weights supplied by the electronic-structure calculation:
Spin-resolved splitting
In noncollinear calculations with spin–orbit coupling (SOC), Bloch states are spinors and lack a definite spin projection. The spin-resolved splitting for the spin component is constructed from pairs of nearby bands via a spin-contrast projection,
where is the band-resolved spin expectation and regularizes accidental zeros in the spin contrast (values as small as are sufficient in practice). For a chosen set of band pairs (e.g., the topmost occupied spinor bands), the effective splitting is the arithmetic mean
In the collinear (no-SOC) case with quantization axis , the splitting is nonzero only for :
where is the set of bands included in the average.
Multipolar moments ()
The spin-resolved energy texture is analyzed through the first three BZ moments [4]:
By parity, and are odd under and diagnose odd-parity () components, whereas is even and diagnoses the even-parity () component. All averages are evaluated on the full 3D mesh.
Strength measures and fractional weights
For each spin channel , the overall strength of -, -, and -wave content is quantified by the norms
The corresponding fractional weights are
and summarize the dominant multipolar character in each spin channel. It is often convenient to report both per-channel fractions and overall fractions [4, 5].